2) There are 2n objects each of three kinds given to two persons, such that each person gets 3n objects. The number of ways of such distribution is equal to


a)   3n2 + 3n + 1


b)   4n-1C3n


c)   3n-1C2n       


d)   3n2 - 2 

Grade:upto college level

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

The answer is option a

Regards
Vikas TU
3 years ago
Therefore, the 3 groups can be chosen 84 x 20 x 1 = 1680 ways. However, since the order of the 3 groups doesn't matter, we have to divide 1680 by 3!. Hence, the number of ways 9 people can be divided into 3 groups is 1680/3! = 1680/6 = 280